OpenStudy (aravindg):

find maximum and minimum values for the function g(x)=x^3-3x

OpenStudy (aravindg):

@hartnn

hartnn (hartnn):

what u got as first derivative ?

OpenStudy (aravindg):

3x^2-3=0 x^2=1 x=1,-1

OpenStudy (aravindg):

should i take second derivative?

hartnn (hartnn):

correct. here u can just put the values in g(x) and by inspection u can tell which is max

hartnn (hartnn):

mathematically, u should take 2nd derivative...

OpenStudy (aravindg):

max at x=-1 and minimum at x=1

OpenStudy (aravindg):

bt then why we took second derivative in last qn?

hartnn (hartnn):

because we can't find the max or min just by putting the points and inspection, can we ?

OpenStudy (aravindg):

i see is my final answers ryt?

OpenStudy (aravindg):

....

OpenStudy (aravindg):

u dr??

hartnn (hartnn):

sorry, got disconnected...yes, its correct

OpenStudy (aravindg):

one last qn and i am done ,,well for this qn i think i gt crct anser bt i want t o knw if i need to take second derivative "find maximum and minimum value of 3x^4-8x^3+12x^2-48x+25

OpenStudy (aravindg):

,,,

hartnn (hartnn):

yes, u need to..

OpenStudy (aravindg):

are answers 0 and 2?

hartnn (hartnn):

mim at x=2

OpenStudy (aravindg):

max at x=0 min at x=2 ryt?

hartnn (hartnn):

yes, both are correct..i had doubt with max at x=0

OpenStudy (aravindg):

thanx a lot!!!! it really helped cant tell u hw thankful i am !!

OpenStudy (aravindg):

seeya gn8 sweeet drms